OpenStudy (anonymous):

If f(x) = (ln x)^2 then f''( sqrt of e) is equal to

OpenStudy (anonymous):

im not sure how to find the first or second derivative

OpenStudy (anonymous):

so are you trying to find the second derivaitve, cause the question is worded weird?

OpenStudy (anonymous):

YES im trying to find the second derivative and then evaluate that at teh square root of e. But you need to find the 1st der to get the 2nd one

OpenStudy (anonymous):

Hint, use the chain rule: \[ \frac{d(ln(u))}{dx} = \frac{1}{u} u ' \]

OpenStudy (anonymous):

But isnt this the derivative of ln(x) to a power not a function

OpenStudy (anonymous):

The same rule applies...

OpenStudy (anonymous):

Only now you will have to use the power rule :-)

OpenStudy (anonymous):

ok so i got the 2nd der. as 2-2log (swrt of e) all divided by (sqrt of e)^2

OpenStudy (anonymous):

what would teh log of sqrt of e be?

OpenStudy (anonymous):

Could you be a bit clearer? The 1st derivative should be: \( 2(ln(x)) \frac{1}{x} \), right?

OpenStudy (anonymous):

yeah